# Chiral superconductivity from parent Chern band and its non-Abelian generalization

### v1

@@ Abstract
We propose a minimal model starting from a parent Chern band with quartic dispersion that can describe the spin-valley polarized electrons in rhombohedral tetra-layer graphene. The interplay between repulsive and attractive interactions on top of that parent Chern band is studied. We conduct standard self-consistent mean-field calculations, and find a rich phase diagram that consists of metal, quantum anomalous Hall crystal, chiral topological superconductor, as well as trivial gapped Bose-Einstein condensate. In particular, there exists a topological phase transition from the chiral superconductor to the Bose-Einstein condensate at zero temperature. Motivated by the recent experimental and theoretical studies of composite Fermi liquid in rhombohedral stacked multi-layer graphene, we further generalize the physical electron model to its composite fermion counterpart based on a field theory analysis. We argue that a chiral spin liquid phase can emerge in the vicinity of the chiral superconductor phase of the composite fermion, i.e., the non-Abelian Moore-Read quantum Hall phase. Our work suggests rhombohedral multi-layer graphene as a potential platform for rich correlated topological phases.
@@
Introduction– Topological superconductor (TSC) is one of the most attractive concepts in modern condensed matter physics. It can host non-Abelian quasiparticles, which are the potential building blocks of quantum computation [1, 2, 3, 4, 5, 6]. In the past decades [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], experimental efforts have been made in searching for chiral topological superconductors, in which the time reversal symmetry is broken. Recent experimental and theoretical progress in graphene-based two-dimensional materials provide a new platform for unconventional superconductors and correlated insulators [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 28, 37, 38, 39, 40, 41, 42]. While the moiré superlattice often plays an important role in those systems [25, 43, 26, 27, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63], it has been shown that the rhombohedral stacked multi-layer graphene exhibits chiral TSC and gate-tunable Chern bands in the absence of moiré pattern, where TSC and quantum anomalous Hall (QAH) insulator are observed at the same electron density [30, 36, 28, 37, 38, 29].
@@
The intertwinement of topology and correlations in graphene-based two-dimensional materials can lead to rich physics. Motivated by this, we propose a minimal model starting from a parent Chern band with quartic dispersion found in rhombohedral tetra-layer graphene [28]. We consider the interplay between repulsive and attractive interactions among electrons. As a concrete example, we choose the Coulomb interaction as the repulsive interaction and the electron-phonon coupling as the attractive one, two of which are the most commonly discussed interactions in solids. Our self-consistent mean-field calculations show a rich phase diagram consists of metal, QAH crystal, chiral TSC, as well as the trivial gapped Bose-Einstein condensate (BEC). In particular, we show there exists a topological phase transition between TSC and BEC phases at zero temperature. Motivated by the recent experimental and theoretical studies of composite Fermi liquid [36, 31, 64, 34, 33, 32, 35], we further provide a field theory analysis that extend our investigation for physical electrons moving under the Berry curvature to composite fermions coupled to dynamical gauge fields, and argue that a chiral spin liquid (CSL) phase [65, 66, 67, 68, 69, 70, 71, 72, 73] can emerge in the vicinity of the chiral TSC phase of the composite fermion [74], i.e., the non-Abelian Moore-Read quantum Hall phase [75]. Our work shed light on investigating varies correlated topological phases in graphene-based materials.
@@
The model– We consider the following minimal model that captures the spin-valley polarized electrons projected to a single parent band,
@@
Here, $H_{0}=\sum_{\bm{k}}\epsilon({\bm{k}})c^{\dagger}_{\bm{k}}c_{\bm{k}}$ is the kinetic term, where $c_{\bm{k}}^{\dagger}$ is the creation operator of an electron with momentum ${\bm{k}}$ in the parent band. According to the spin-valley polarization, the electron is single-flavored. In rhombohedral tetra-layer graphene, the parent band is believed to be a Chern band with quartic dispersion $\epsilon({\bm{k}})=\alpha|{\bm{k}}|^{4}$ [28], and the Brillouin zone (BZ) is chosen to be hexagonal to respect the $C_{3}$ symmetry. The electron filling is fixed to $\nu=1$. The interaction $H_{\rm int}$ can be divided into the repulsive part and the attractive part. As an example, we consider the repulsive interaction from the electron-electron Coulomb interaction $H_{\text{e-e}}$ and the attractive interaction from the electron-phonon coupling $H_{\text{e-ph}}$ [76], both of which are projected to the parent Chern band,
@@
with $\tilde{V}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=vV({{\bm{k}}_{1}-{\bm{k}}_{4}}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$ and $\tilde{U}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=uU({{\bm{k}}_{1}-{\bm{k}}_{4}}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$. Here $u,v$ are positive coupling constants that denote the interaction strengths. The electron-electron interaction is given by the gate-screened Coulomb potential $V({\bm{q}})=\tanh{d|{\bm{q}}|}/(d|{\bm{q}}|)$, while the electron-phonon interaction takes the form of $U({\bm{q}})=q_{0}^{2}/(|{\bm{q}}|^{2}+q_{0}^{2})$. In this work, we fix $d=1.0/|{\bm{K}}|$ around the order of the screening length [31, 76], and $q_{0}=0.3|{\bm{K}}|$ satisfying $q_{0}<|{\bm{K}}|$ [77] [^1] In numerical calculations we find the superconducting phase is relatively inert to $d$, but rather sensitive to $q_{0}$ which is consistent with Ref. [77]., where ${\bm{K}}$ denotes the momentum at the $K$-point of the hexagonal BZ. We then define a energy scale for kinetic term $H_{0}$ as $\epsilon_{K}=\alpha|{\bm{K}}|^{4}$, and evaluate $(u,v)$ in terms of $\epsilon_{K}$. The projected form factor ${\mathcal{F}}$ encodes the quantum geometry of the parent band [76],
@@
Self-consistent calculation– The competition between $H_{\text{e-e}}$ and $H_{\text{e-ph}}$ motivates us to investigate the full phase diagram in terms of $u$ and $v$. To see this, we conduct self-consistent mean-field calculations for the model Hamiltonian Eq. (1). We start from both particle-particle and particle-hole mean-field channels, and seek for self-consistent solutions for different $(u,v)$. For simplicity, we define $S({\bm{q}})=vV({\bm{q}})-uU({\bm{q}})$, such that:
@@
For the particle-particle channel, we decompose the electron momentum as $({\bm{k}}+{\bm{g}})$, where ${\bm{k}}$ is defined in the first BZ, and ${\bm{g}}_{i}$ labels the reciprocal lattice vectors under $C_{3}$ rotation. Under the Hatree-Fock approximation [76], the self-consistent mean-field interaction Hamiltonian can be divided into the Hartree term with ${\bm{k}}_{1}={\bm{k}}_{4}$ and ${\bm{k}}_{2}={\bm{k}}_{3}$,
@@
Figure 1: (1) Phase diagram of Hamiltonian Eq. (1) from self-consistent calculations. The gray corner shows the possible metallic phase, to the accuracy of our numerics. The yellow block stands for the QAH crystal phase. The blue region denotes the $(p+ip)$-wave TSC, and the purple block represents the trivial gapped BEC phase. (b-d) Schematic phase diagrams for the composite fermion theory Eq. (9), with (b) small Maxwell coupling $\kappa$, (c) negative composite fermion mass $m$, (d) large Maxwell coupling $\kappa$.
For the particle-hole channel, we assume that the pairing happens between electrons with opposite momenta. In this way, the BdG Hamiltonian reads,
@@
where $\Psi_{\bm{k}}=(c_{\bm{k}},c^{\dagger}_{-{\bm{k}}})^{T}$ is the Nambu spinor. Here $\mu$ is the chemical potential determined by the self-consistent equation of electron number $N$ at filling $\nu=1$. When $\mu<0$, the BdG Hamiltonian Eq. 7 describes a BEC phase. The superconducting gap
$\Delta({\bm{k}})=\sum_{{\bm{k}}^{\prime}}\frac{1}{2}[S({\bm{k}}-{\bm{k}}^{\prime}){\mathcal{F}}_{{\bm{k}},{\bm{k}}^{\prime}}^{2}-S({\bm{k}}+{\bm{k}}^{\prime}){\mathcal{F}}_{-{\bm{k}},{\bm{k}}^{\prime}}^{2}\langle c_{-{\bm{k}}^{\prime}}c_{{\bm{k}}^{\prime}}\rangle,$ (8)
@@
respects spatial asymmetry $\Delta(-{\bm{k}})=-\Delta({\bm{k}})$ representing the odd parity self-consistent solution of Eq. (7) for the pairing between single-flavored electrons. We also denote $E({\bm{k}})=\sqrt{[\epsilon({\bm{k}})-\mu]^{2}/4+|\Delta({\bm{k}})|^{2}}$ as the dispersion of the Bogoliubov band.
@@
Phase diagram– With above, we conduct self-consistent calculations of Eq. (1) for the particle-particle and the particle-hole channels numerically. We check the convergence of these two channels for each pair of $(u,v)$, and the phase diagram Fig. [1.(a)] is mapped out in terms of $u$ and $v$ varying from 00 to $\epsilon_{K}$. If both channels converge, we pick the energetic favored one. We find that, when both $u$ and $v$ are small ($u,v<0.05\epsilon_{K}$), the system favors a metallic phase, as shown in the gray region. A strong repulsive Coulomb interaction (i.e., large $v$) can drive the system to an insulating phase, with a finite charge gap $\delta$ open at the first BZ edge, as shown in the yellow region. Within the yellow region, the valence band of $H_{\rm HF}$ has non-trivial Chern number and order parameter $P_{{\bm{g}}_{1}{\bm{g}}_{2}}({\bm{k}})$, indicating a QAH crystal phase. Note that, deep in the QAH crystal region, the band is relative flat and the distribution of Berry curvature is rather homogeneous within the first BZ, as shown in the Fig. [2]. We conjecture that the valence band is very close to the lowest Landau level which can give rise to fractional Chern insulators in the presence of moiré potential at fractional $\nu$ [53, 54, 33, 32, 34, 35, 31].
Figure 2: The ratio between band width $w$ and band gap $\delta$ with respect to the interacting strength $u$ and $v$ within the region of QAH crystal phase. The blue, orange, and green curve is plotted at $u=0.2\epsilon_{K}$, $u=0.6\epsilon_{K}$, and $u=1.0\epsilon_{K}$, respectively. One can see that the band width $w$ is small (less than 0.1) compared with the charge gap $\delta$ for all the $(u,v)$ we consider. The inset shows the distribution of Berry curvature within first BZ for $u=0$, $v=0.2\epsilon_{K}$, whose integral gives Chern number $C=1$.
@@
For the $(u,v)$ in the blue region of Fig. [1.(a)], one can find energetic favored convergent solutions for BdG Hamiltonian Eq. (7). The superconducting gap distribution within the first BZ is shown in Fig. [3], whose real (imaginary) part has the profile of $p_{x}$-wave ($p_{y}$-wave) pairing, indicating a chiral TSC phase with $(p+ip)$-wave pairing symmetry [5]. The physical origin for the chiral TSC can be understood as following. For a trivial parent band, due to the Cooper instability, a finite electron-phonon coupling will lead to $s$-wave superconducting phase. For a parent Chern band, the non-trivial Berry curvature will give rise to $p$-wave superconducting phase. This is in accordance with previous work [77, 79]. The vortices in bulk of this phase can trap non-Abelian Majorana zero modes [5].
Figure 3: Superconducting gap distribution within the first BZ with $u=0.1\epsilon_{K}$, $v=0.05\epsilon_{K}$. The dashed line denotes the fermi surface at $\nu=1$. The left and right panel denotes, respectively, the real and imaginary part of $\Delta({\bm{k}})$, which clearly shows the feature of a $(p+ip)$-wave superconductor.
@@
In the purple region of Fig. [1.(a)] we have $\mu<0$ that stands for the BEC phase. The transition between TSC and BEC phases here is rather interesting. The TSC phase has positive chemical potential $\mu$. Deep in the TSC phase, the electrons are dominantly paired close to the Fermi surface, in accordance with the fact that the Bogoliubov band gap is approximately minimized near the first BZ edge deep in the TSC phase. This can be checked for the inset of $u=0.1\epsilon_{K}$ in Fig. [4] showing the corresponding Bogoliubov band dispersion. The Cooper pair is extended in real space. The increasing of electron-phonon interaction $u$ leads to the drop of $\mu$ as well as the decreasing of the Cooper pair size, such that the electrons deep in the Fermi sea starts to pair. Different from the case of $s$-wave, for $(p+ip)$-wave there exists a critical $u^{*}$ where $\mu=0$ and Bogoliubov band gap closes at the $\Gamma$ point. This can be seen that $E({\bm{k}}=0)=0$ when $\mu=0$ and $\Delta({\bm{k}}=0)=0$ for $(p+ip)$-wave. The Chern number for the Bogoliubov band changes from $1$ in TSC phase to 00 in BEC phase, signaling a topological phase transition from the TSC to the trivial gapped BEC phase. In the BEC phase, the more negative $\mu$ is, the more electrons in the Fermi sea are paired and the size of the Cooper pairs are smaller. The gap of Bogoliubov band is minimal at $\Gamma$ point, as shown in the insets with $u=0.47\epsilon_{K}$ and $u=0.6\epsilon_{K}$ in Fig. [4]. Details of the topological phase transition from TSC to BEC phases with fixed $v=0.05\epsilon_{K}$ and varying $u$ are shown in Fig. [4]. We point out that at finite temperature such a phase transition will change to crossover [80].
@@
Figure 4: Transition between the chiral TSC (blue region) and the trivial BEC phase (purple region), with $v=0.05\epsilon_{K}$ and $u\in[0.1,0.6]\epsilon_{K}$. One can see the chemical potential ($\mu$) drop as the increasing of the electron-phonon coupling strength $(u)$. The bottom insets illustrate the band structure of Bogoliubov quasiparticles. The band gap closes at the critical point $(u^{*}\approx 0.33\epsilon_{K},\mu=0)$, indicating the topological phase transition from chiral TSC to trivial BEC phase.
@@
Generalization to composite fermion theory– The phase diagram Fig. [1.(a)] provides insights to study the potential non-Abelian quantum Hall phase in rhomobohedral multi-layer graphene, in which the composite fermi liquid has been studied both experimentally [36] and theoretically [31]. It is natural to extend our model Eq. (1) with physical electrons moving in the background Berry curvature to that of the composite fermion coupled to the dynamical gauge field. The chiral TSC phase for the composite fermion is proposed to be described by the Moore-Read [75, 1] quantum Hall state, which exhibits non-Abelian topological order. Below we generalize the phase diagram Fig. [1.(a)] into the composite fermion case with the help of field theory. The composite fermions in the presence of pairing can be modeled as [74, 81, 82]
@@
$+$ $|D_{2a}\Phi|^{2}+r|\Phi|^{2}+t|\Phi|^{4}-\lambda\left(\Phi^{\dagger}\psi^{\beta}\psi_{\beta}+h.c.\right)$
-- $\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A,$
where $m$ is the Dirac mass of the composite fermion $\psi$, which is charged under the dynamical compact $U(1)$ gauge field $a$, and $\kappa$ is the Maxwell coupling of $a$. Coupling constants $t$ and $\lambda$ are set to be positive. Dynamical compact $U(1)$ gauge field $b$ is coupled to the background gauge field $A$ through a mutual Chern-Simons (CS) term, whose flux corresponds to the current of the physical electron. The Maxwell coupling of $b$ is neglected since it is more irrelevant than the self CS term. The complex boson field $\Phi$ describes the pairing of the composite fermion, which is consequently charge-2 under the gauge field $a$. The $\lambda$-term describes $s$-wave superconductivity for the Dirac composite fermion, which is equivalent to the chiral TSC for the non-relativistic composite fermion [74]. This is due to the intrinsic Berry curvature of the Dirac fermion, with a similar mechanism that Berry curvature can turn the $s$-wave superconductor into a chiral $(p+ip)$-wave one in the physical fermion case [77, 83].
The phase diagram of Eq. (9) can be analyzed at the mean-field level. We first consider the non-superconducting phases. This corresponds to $r>0$, where $\Phi$ is gapped and $\psi$ can be integrated out. The resulting topological part of the Lagragian reads

### v3

# Chiral superconductivity from parent Chern band and its nonabelian generalization

@@ Abstract
We propose a minimal model starting from a parent Chern band with quartic dispersion that can describe the spin-valley polarized electrons in rhombohedral tetralayer graphene. The interplay between repulsive and attractive interactions on top of that parent Chern band is studied. We conduct standard self-consistent mean-field calculations, and find a rich phase diagram that consists of metal, quantum anomalous Hall crystal, chiral topological superconductor, as well as trivial gapped Bose–Einstein condensate. In particular, there exists a topological phase transition from the chiral superconductor to the Bose–Einstein condensate at zero temperature. Motivated by the recent experimental and theoretical studies of composite Fermi liquid in rhombohedral stacked multilayer graphene, we further generalize the physical electron model to its composite fermion counterpart based on a field theory analysis. The chiral superconductor phase of the composite fermion becomes the nonabelian Moore–Read quantum Hall phase. We argue that a chiral (pseudo-)spin liquid phase can emerge in the vicinity of this Moore–Read quantum Hall phase. Our work suggests rhombohedral multilayer graphene as a potential platform for rich correlated topological phases.
@@
Introduction— Topological superconductor (TSC) is one of the most attractive concepts in modern condensed matter physics. It can host nonabelian quasiparticles, which are the potential building blocks of quantum computation [1, 2, 3, 4, 5, 6]. In the past decades [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], experimental efforts have been made in searching for chiral topological superconductors, in which the time reversal symmetry is broken. Recent experimental and theoretical progress in graphene-based two-dimensional materials provide a new platform for unconventional superconductors and correlated insulators [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 28, 37, 38, 39, 40, 41, 42, 43]. While the moiré superlattice often plays an important role in those systems [25, 44, 26, 27, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64], it has been shown that the rhombohedral stacked multilayer graphene exhibits chiral TSC and gate-tunable Chern bands in the absence of moiré pattern, where TSC and anomalous Hall crystal (AHC) are observed at the same electron density [30, 36, 28, 37, 38, 29].
@@
The intertwinement of topology and correlations in graphene-based two-dimensional materials can lead to rich physics. Motivated by this, we propose a minimal model starting from a parent Chern band with quartic dispersion found in rhombohedral tetralayer graphene [28]. We consider the interplay between repulsive and attractive interactions among electrons. As a concrete example, we choose the Coulomb interaction as the repulsive interaction and the electron-phonon coupling as the attractive one, two of which are the most commonly discussed interactions in solids. Note that the formation of the anomalous Hall crystal has been proposed in Ref. [65], breaking the translation symmetry explicitly and providing the possibility for electron-phonon coupling. Our self-consistent mean-field calculations show a rich phase diagram consisted of metal, AHC, chiral TSC, as well as the trivial gapped Bose-Einstein condensate (BEC). In particular, we show there exists a topological phase transition between TSC and BEC phases at zero temperature. Motivated by the recent experimental and theoretical studies of composite Fermi liquid [36, 31, 66, 34, 33, 32, 35], we further provide a field theory analysis that extend our investigation for physical electrons moving under the Berry curvature to composite fermions coupled to dynamical gauge fields, and argue that a chiral (pseudo-)spin liquid (CSL) phase [67, 68, 69, 70, 71, 72, 73, 74, 75] can emerge in the vicinity of the chiral TSC phase of the composite fermion [76], i.e., the nonabelian Moore–Read quantum Hall phase [77]. Our work sheds light on investigating various correlated topological phases in graphene-based materials.
@@
The model— We consider the following minimal model that captures the spin-valley polarized electrons projected to a single parent band,
@@
Here, $H_{0}=\sum_{\bm{k}}\epsilon({\bm{k}})c^{\dagger}_{\bm{k}}c_{\bm{k}}$ is the kinetic term, where $c_{\bm{k}}^{\dagger}$ is the creation operator of an electron with momentum $\bm{k}$} in the parent band. According to the spin-valley polarization, the electron is single-flavored. In rhombohedral tetralayer graphene, the parent band is believed to be a Chern band with quartic dispersion $\epsilon({\bm{k}})=\alpha|{\bm{k}}|^{4}$ [28], and the Brillouin zone (BZ) is chosen to be hexagonal to respect the $C_{3}$ symmetry. The electron filling is fixed to $\nu=1$. The interaction $H_{\rm int}$ can be divided into the repulsive part and the attractive part. As an example, we consider the repulsive interaction from the electron-electron Coulomb interaction $H_{\text{e-e}}$ and the attractive interaction from the electron-phonon coupling $H_{\text{e-ph}}$ [78], both of which are projected to the parent Chern band,
@@
with $\tilde{V}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=vV({{\bm{k}}_{1}-{\bm{k}}_{4}}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$ and $\tilde{U}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=uU({{\bm{k}}_{1}-{\bm{k}}_{4}}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$. Here $u,v$ are positive coupling constants that denote the interaction strengths. The electron-electron interaction is given by the gate-screened Coulomb potential $V({\bm{q}})=\tanh{d|{\bm{q}}|}/(d|{\bm{q}}|)$, while the electron-phonon interaction takes the form of $U({\bm{q}})=q_{0}^{2}/(|{\bm{q}}|^{2}+q_{0}^{2})$  [^1] Note that, this form of electron-phonon interaction asymptotically approaches the delta function as $q_{0}\rightarrow 0$. See Ref. [97] for further details.. In this work, we fix $d=10.0/|{\bm{K}}|$ at the order of the screening length [31, 78], and $q_{0}=0.3|{\bm{K}}|$ satisfying $q_{0}<|{\bm{K}}|$ [80] [^2] In numerical calculations we find the superconducting phase is relatively inert to $d$, but rather sensitive to $q_{0}$ which is consistent with Ref. [80]. When $q_{0}$ is enlarged, the area of BEC phase grows, while the area of TSC phase shrinks. This suggests that a more extended electron-phonon coupling will favor superconductivity but suppress topological physics., where $\bm{K}$} denotes the momentum between the nearest two inequivalent $K$-points of the hexagonal BZ. We then define the energy scale for kinetic term $H_{0}$ as $\epsilon_{K}=\alpha|{\bm{K}}|^{4}$, and evaluate $(u,v)$ in terms of $\epsilon_{K}$. The projected form factor $\mathcal{F}$} encodes the quantum geometry of the parent band [78],
@@
Self-consistent calculation— The competition between $H_{\text{e-e}}$ and $H_{\text{e-ph}}$ motivates us to investigate the full phase diagram in terms of $u$ and $v$. To see this, we conduct self-consistent mean-field calculations for the model Hamiltonian Eq. (1). We start from both particle-particle and particle-hole mean-field channels, and seek for self-consistent solutions for different $(u,v)$. For simplicity, we define $S({\bm{q}})=vV({\bm{q}})-uU({\bm{q}})$, such that:
@@
For the particle-hole channel, we decompose the electron momentum as $({\bm{k}}+{\bm{g}}_{i})$, where $\bm{k}$} is defined in the first BZ, and ${\bm{g}}_{i}$ labels the reciprocal lattice vectors under $C_{3}$ rotation. Under the Hartree–Fock approximation [78], the self-consistent mean-field Hamiltonian can be divided into the Hartree term with ${\bm{k}}_{1}={\bm{k}}_{4}$ and ${\bm{k}}_{2}={\bm{k}}_{3}$,
@@
Figure 1: (a) Phase diagram of Hamiltonian Eq. (1) from self-consistent calculations. The gray region shows the (semi-)metallic phase. The yellow block stands for the AHC phase. The purple paddles denote the $(p+ip)$-wave TSC, and the blue block represents the trivial gapped BEC phase. The inset zooms in the left bottom corner of the phase diagram with higher resolution.
(b-d) Schematic phase diagrams for the composite fermion theory Eq. (8) under fixed $\lambda,t>0$, with (b) small Maxwell coupling $\kappa$, (c) negative composite fermion mass $m$, (d) large Maxwell coupling $\kappa$.
For the particle-particle channel, we assume that the pairing happens between electrons with opposite momenta. In this way, the BdG Hamiltonian reads,
@@
where $\Psi_{\bm{k}}=(c_{\bm{k}},c^{\dagger}_{-{\bm{k}}})^{T}$ is the Nambu spinor. Here $\mu$ is the chemical potential determined by the self-consistent equation of electron number NN at filling $\nu=1$. When $\mu<0$, the BdG Hamiltonian Eq. 7 describes a BEC phase. The superconducting gap which respects spatial asymmetry $\Delta(-{\bm{k}})=-\Delta({\bm{k}})$ representing the odd parity self-consistent solution of Eq. (7) for the pairing between single-flavored electrons. We also denote $E({\bm{k}})=\sqrt{[\epsilon({\bm{k}})-\mu]^{2}/4+|\Delta({\bm{k}})|^{2}}$ as the dispersion of the Bogoliubov band.
@@
Phase diagram— With above, we conduct self-consistent calculations of Eq. (1) for the particle-particle and the particle-hole channels numerically and map out the phase diagram Fig. [1.(a)] with $0\leq u\leq 5\epsilon_{K}$ and $0\leq v\leq 10\epsilon_{K}$. We find that, when the values of $u$ and $v$ are approximately comparable, the system favors a (semi-)metallic phase, as shown in the gray region. A strong repulsive Coulomb interaction (i.e., large $v$) can drive the system to an insulating phase, with a finite charge gap open at the first BZ edge, as shown in the yellow region. Within the yellow region, the valence band of $H_{\rm HF}$ has nontrivial Chern number $C=1$ and order parameter $P_{{\bm{g}}_{1}{\bm{g}}_{2}}({\bm{k}})$, indicating an AHC phase.
@@
For the $(u,v)$ in the purple region of Fig. [1.(a)], one can find energetically favored convergent solutions for BdG Hamiltonian Eq. (7). The superconducting gap distribution is shown in the inset of Fig. [2], whose real (imaginary) part has the profile of $p_{x}$-wave ($p_{y}$-wave) pairing, indicating a chiral TSC phase with $(p+ip)$-wave pairing symmetry [5] [^3] For the $C_{3}$ symmetric system, the $p$-wave and the $d$-wave superconductors are related by a time reversal transformation, since the former one has angular momentum $l=+1$ and the latter has l=+2=1l=+2=-1 (mod 3).. The physical origin for the chiral TSC can be understood as follows. For a trivial parent band, due to the Cooper instability, a finite electron-phonon coupling will lead to $s$-wave superconducting phase. For a parent Chern band, the nontrivial Berry curvature will give rise to $p$-wave superconducting phase. This is in accordance with previous work [80, 83, 84]. The vortices in the bulk of this phase can trap nonabelian Majorana zero modes [5].
@@
In the blue region of Fig. [1.(a)] we have $\mu<0$ that stands for the BEC phase. The transition between TSC and BEC phases here is rather interesting. The TSC phase has positive chemical potential $\mu$. Deep in the TSC phase, the electrons are dominantly paired close to the Fermi surface, in accordance with the fact that the Bogoliubov band gap is approximately minimized near the Fermi wavevector deep in the TSC phase. This can be checked for the inset of $u=0.72\epsilon_{K}$ in Fig. [2] showing the corresponding Bogoliubov band dispersion. The Cooper pair is extended in real space. The increasing of electron-phonon interaction $u$ leads to the drop of $\mu$ as well as the decreasing of the Cooper pair size, such that the electrons deep in the Fermi sea start to pair. Different from the case of $s$-wave, for $(p+ip)$-wave there exists a critical $u^{*}$ where $\mu=0$ and Bogoliubov band gap closes at ${\bm{k}}=0$. This can be seen that $E({\bm{k}}=0)=0$ when $\mu=0$ and $\Delta({\bm{k}}=0)=0$ for $(p+ip)$-wave. The Chern number for the Bogoliubov band changes from 11 in TSC phase to 0 in BEC phase, signaling a topological phase transition from the TSC to the trivial gapped BEC phase. In the BEC phase, the more negative $\mu$ is, the more electrons in the Fermi sea are paired and the size of the Cooper pairs is smaller. The gap of Bogoliubov band is minimal at ${\bm{k}}=0$, as shown in the insets with $u=1.08\epsilon_{K}$ and $u=1.2\epsilon_{K}$ in Fig. [2]. Details of the topological phase transition from TSC to BEC phases with fixed $v=\epsilon_{K}$ and varying $u$ are shown in Fig. [2]. We point out that at finite temperature such a phase transition will change to a crossover [85].
@@
Figure 2: Transition between the chiral TSC (purple region) and the trivial BEC phase (blue region), with $v=\epsilon_{K}$ and $0.7\epsilon_{K}\leq u\leq 1.2\epsilon_{K}$. One can see the chemical potential ($\mu$) drop as the increasing of the electron-phonon coupling strength $(u)$. The inset in TSC regime shows superconducting gap distribution with $u=0.72\epsilon_{K}$, $v=\epsilon_{K}$, where the real and imaginary part of $\Delta({\bm{k}})$ is shown in the left and right panel, clearly suggesting the feature of a $(p+ip)$-wave superconductor. The bottom insets illustrate the band structure of Bogoliubov quasiparticles. The band gap closes at the critical point $(u^{*}\approx 0.97\epsilon_{K},~\mu=0)$, indicating the topological phase transition from chiral TSC to trivial BEC phase.
@@
Generalization to composite fermion theory— The phase diagram Fig. [1.(a)] provides insights to study the potential nonabelian quantum Hall phase in rhombohedral multilayer graphene, in which the composite Fermi liquid has been studied both experimentally [36] and theoretically [31]. It is natural to extend our model Eq. (1) with physical electrons moving in the background Berry curvature to that of the composite fermion coupled to the dynamical gauge field. The chiral TSC phase for the composite fermion is proposed to be described by the Moore–Read [77, 1] quantum Hall state, which exhibits nonabelian topological order. Below we generalize the phase diagram Fig. [1.(a)] into the composite fermion case with the help of field theory. The composite fermions in the presence of pairing can be modeled as [76, 86, 87]
@@
$+|D_{2a}\Phi|^{2}+r|\Phi|^{2}+t|\Phi|^{4}-\lambda\left(\Phi^{\dagger}\psi_{+}\psi_{-}+h.c.\right)$
$-\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A.$
Here gamma matrices can be chosen as $\gamma^{0}=\sigma^{z}$, $\gamma^{1}=i\sigma^{y}$, and $\gamma^{2}=-i\sigma^{x}$ [76]. The composite fermion $\psi$ is a two-component Dirac fermion with mass $m$. It is charged under the dynamical compact $U(1)$ gauge field $a$, and $\kappa$ is the Maxwell coupling of $a$. The complex boson field $\Phi$ describes the pairing of the composite fermion, which is consequently charge-2 under the gauge field $a$. Its condensation is controlled by $r$, with the coupling constant of the quartic term $t$ set to be large and positive. The $\lambda$-term ($\lambda>0$) describes $s$-wave pairing for the Dirac composite fermion, which is equivalent to the chiral TSC for the nonrelativistic composite fermion [76]. This is due to the intrinsic Berry curvature of the Dirac fermion, with a similar mechanism that Berry curvature can turn the $s$-wave superconductor into a chiral one in the physical fermion case [80, 88]. Dynamical compact $U(1)$ gauge field $b$ is coupled to the background gauge field AA through a mutual Chern–Simons (CS) term, whose flux corresponds to the current of the physical electron. The Maxwell coupling of $b$ is neglected since it is more irrelevant than the self CS term.
The phase diagram of Eq. (8) can be analyzed at the mean-field level. We first consider the nonpairing phases. This corresponds to $r>0$, where $\Phi$ is gapped and $\psi$ can be integrated out. The resulting topological part of the Lagrangian reads